MÜNCHENER STOCHASTIK-TAGE 1998 Sekt.-Nr.:12

Stochastische (partielle) Differentialgleichungen

Sektionsleiter: H. Engelbert, Jena

Breckner , Hannelore Martin-Luther Universität Halle-Wittenberg

Steuerung der Stochastischen Navier-Stokes Gleichung

Es wird ein Steuerungsproblem für die stochastische Navier-Stokes Gleichung bezüglich der äußeren Kraft betrachtet. Diese Gleichung wird als Itogleichung über einem Evolutionstripel von Hilberträumen (V, H, V^{∗}) interpretiert

u(t) + ∫_{0}^{t} Au(s)ds = u_{0} + ∫_{0}^{t} B(u(s), u(s))ds + ∫_{0}^{t} f(s)ds + ∫_{0}^{t} Cu(s)dw(s)  (1)

für alle t ∈ [0, T] und fast alle ω ∈ Ω. Es gelten folgende Voraussetzungen:

• (Ω, F, P) ist ein vollständiger Wahrscheinlichkeitsraum, {F_{t} | t ∈ [0, T]} ist eine Filtration bezüglich eines reellen Wiener Prozeßes {w_{t} | t ∈ [0, T]};

• A : D(A) ⊆ H → H ist ein linearer, selbstadjungierter, koerzitiver Operator;

• f : [0, T] × Ω → H ist eine {F_{t} | t ∈ [0, T]} - meßbare Funktion, deren Norm fast überall beschränkt ist;

• B : V × V → V^{∗} ist eine bilineare Abbildung mit

|⟨B(u, v), z⟩| ≤ β∥z∥_{V} ∥u∥_{V}^{1/2}∥u∥_{H}^{1/2}∥v∥_{V}^{1/2}∥v∥_{H}^{1/2}

für jedes u, v, z ∈ V (β ist eine positive Konstante) und ⟨B(u, v), v⟩ = 0 für alle u, v ∈ V;

• C : H → H ist eine lineare und stetige Abbildung;

• u_{0} ∈ L_{H}^{2}(Ω) ist F_{0}-meßbar. Es wird folgendes stochastisches Steuerungsproblem betrachtet:

J(f) = E ∫_{0}^{T} ∥f(x, t)∥^{2}dt + E ∫_{0}^{T} ∥e^{−c ∫_{0}^{t} ∥u_{f} (s)∥_{V}^{2} ds}u_{f}(t)∥_{V}^{2} dt ⟶ min_{f}

wobei u_{f} die Lösung von (1) ist, wenn die Kraft f wirkt (c > 0 ist ein Diskontierungsfaktor). Es wird die Existenz einer optimalen Steuerung bewiesen und ein stochastisches Maximumprinzip hergeleitet. Insbesondere werden Approximationsmöglichkeiten und Existenzaussagen für den adjungierten Prozeß angegeben.

Buckdahn, Rainer Hauptvortrag Université de Bretagne Occidentale, Brest

Backward Stochastic Differential Equations and Applications

Backward stochastic differential equations (in short BSDE’s) are a new type of stochastic differential equations: Their terminal condition is a given random variable which can depend on the whole path of the underlying Brownien motion while the solution one looks for is a couple of processes adapted to this underlying Brownian motion. The notion of the BSDE generalizes the martingale representation property on the Wiener- and on the Wiener-Poisson space, respectively; see [Pardoux/Peng, 1990],[Buckdahn/Pardoux,1994]. It turns out that the BSDE’s generalize the Feynman-Kac formula: They provide a probabilistic formula for the solution of a certain class of systems of quasi-linear parabolic PDE’s of second order. In particular, the BSDE’s are naturally associated to the viscosity solution of PDE’s [Pardoux/Peng,1992], and the BSDE’s with jumps are naturally associated to the viscosity solution of integral-partial differential equations [Barles/Buckdahn/Pardoux, 1997]. This explains the multitude of applications for BSDE’s. My talk will give a survey about the notion of the BSDE and present some of its apllications on which I have worked, namely its applications i) to a homogenization problem for semi-linear PDE’s [Buckdahn/Hu/Peng, 1996], [Buckdahn/Hu/Peng, 1997]; ii) in viability theory [Buckdahn/Quincampoix/Rascanu, 1997]; iii)viscosity solution of integral-partial differential equations [Barles/Buckdahn/Pardoux, 1997]; iv) in finance: evaluation of contingent claims under constraints in an incomplet financial market [Buckdahn/Hu, 1996].

Engelbert, Hans-Jürgen; Wolf, Jochen Friedrich-Schiller-Universität Jena, Institut für Stochastik

Eine Bemerkung Über den Satz von Nakao

Wir betrachten die eindimensionale stochastische Differentialgleichung

X_{t} = X_{0} + ∫_{0}^{t} a(X_{s}) ds + ∫_{0}^{t} b(X_{s}) ds, t ≥ 0,  (1)

mit einer Brownschen Bewegung (B, IF) und meßbaren reellen Drift- bzw. Diffusionskoeffizienten a und b. S. Nakao (1972) hat für beschränkte a und b die pfadweise Eindeutigkeit der Lösung von (1) gezeigt, wenn

b(x) ≥ c > 0, x ∈ IR,  (2)

gilt und wenn b von endlicher Variation ist. (Eine Übertragung des Ergebnisses auf den Fall, daß die oben formulierten Bedingungen an a und b nur lokal erfüllt sind, erfolgt ohne Schwierigkeit.) In diesem Vortrag geben wir einen einfachen und natürlichen Beweis des formulierten Resultats. Dabei stellt sich heraus, daß die (lokal) endliche Variation von b eine unnatürliche Bedingung ist. Diese liefert automatisch, aber unerwünscht, die (lokale) Beschränktheit von b. Vielmehr ist die (lokal) endliche Variation von b^{−1} von Bedeutung. Für nichtnegative b sichert das bereits (lokal) Bedingung (2). Andererseits existieren unter dieser Voraussetzung auch Koeffizienten b, die nicht (lokal) beschränkt und deshalb auch nicht selbst (lokal) von endlicher Variation sind. (Solche Funktionen müssen auch nicht Lipschitz-stetig sein.) Zusammenfassend erhalten wir Existenz und pfadweise Eindeutigkeit einer (i.a. explodierenden) Lösung von Gleichung (1), wenn nur a lokal integrierbar und b^{−1} nichtnegativ und lokal von endlicher Variation ist. Das Ergebnis bleibt auch gültig für Gleichungen mit verallgemeinerter Drift, wobei dann die Bedingung an a entfällt.

Gawinecki, Jerzy Institute of Mathematics and Operations Research, Military University of Technology

Stochastic thermoelasticity equation

Abstract. We consider the stochastic equation describing thermoelastic materials. It means that we consider hyperbolic-parabolic system of equations in three dimensional space putting on the right hand side of this equations the term \frac{∂W (t,x)}{∂t} which is a meak desivative of Wiener process. The initial value u^{0} is a random variable. Applying the method of Hilbert space we obtain the behaviour of the probability distribution μ_{t} of u(t, ·) where u(t, ·) is the solution of our system.

Grecksch, Wilfried Martin-Luther-Universität Halle-Wittenberg

Eine stochastische partielle Differentialgleichung erster Ordnung

Für die nichtlineare Aufgabe

dX(t, x) = ∑_{i=1}^{n} \frac{∂}{∂x_{i}} (b(t, x)X(t, x))dt + A(t, x, X(t, x))dt + B(t, x, X(t, x))dw(t)  (1)

mit t ∈]0, T], x ∈ R^{n} und X(0, x) = X_{0}(x) wird die Existenz und Eindeutigkeit einer W_{2}^{1}(R^{n})- Lösung und einer verallgemeinerten Lösung bewiesen. Dabei wird die Methode der parabolischen Regularisierung angewendet: Dem Problem (1) werden für ε > 0 über dem Evolutionstripel

(W_{2}^{1}(R^{n}) , L^{2}(R^{n}) , W^{−1}(R^{n}))

die parabolischen Itogleichungen

dX_{ε}(t, x) = εΔ_{x}X_{ε}(t, x)dt + ∑_{i=1}^{n} \frac{∂}{∂x_{i}} (b(t, x)X(t, x))dt

+A(t, x, X(t, x))dt + B(t, x, X(t, x))dw(t)  (2)

mit t ∈]0, T], x ∈ R^{m} und X(0, x) = X_{0}(x) zugeordnet. Unter Glattheitsvoraussetzungen an b, A, B gilt für die Lösungen von (1) und (2): Theorem lim_{ε→0} E sup_{t∈[0,T]} ∫_{R^{n}}(X_{ε}(t, x) − X(t, x))^{2}dx = 0.

Gwinner , Joachim UniBW München, Fak. LRT

Stochastic partial differential equations of elliptic type with unilateral boundary conditions

In porous media flow one often encounters irregular data of the hydraulic conductivity which can be modeled as a stochastic process. Another important feature is the occurence of unilateral boundary conditions, like the Signorini condition, in addition to classical linear boundary conditions. Thus we are led to the study of unilateral stochastic boundary value problems. First in this contribution, we address analytical issues (solution space, existence). Then we turn to numerical solution procedures and investigate their convergence behaviour. The numerical methods considered are stochastic (Monte Carlo methods), deterministic (nonconform finite element methods for spatial discretization), and combinations thereof.

Holban, Evelyna UniBw München, Infomatik

Long time behavior of a capillary barrier

A capillary barrier, on top of a landfill can be a cheap, long-living and effective alternative to a conventional cover. An important requirement of a landfill cover is to prevent the infiltration of the precipitation through the wastes. We model the water flow in layered soils by the use of an elliptic-parabolic partial differential equation. In order to analyze the performance and the reliability of capillary barriers we investigate the behavior of our model during a long time period simulation using precipitation rates generated on the basis of a stochastic daily precipitation model.

Kelanemer, Youcef Universität der Bundeswehr München

Optimal Control of Soil Venting

We consider the optimal control of the problem of soil venting which is governed by two stochastic partial differential equations for the air and pollutant transport in a contaminated site. Our optimization problem is to maximize the quantity of extracted chemicals for given total pumping rate. The transmissivity of the air in the soil and the pollutant volatilization are stochastic fields. The control variable is the discharges of some wells. Some theoretical, numerical and computational aspects are considered.

Keller, Hannes Institut für dynamische Systeme, Universität Bremen

Attractors for stochastic differential equations via transformation into random differential equations

Nonautonomous and random dynamical systems and attractors for such systems are introduced. Existence of attractors for random dynamical systems generated by stochastic differential equations of particular structure is proved. For this reason we transform the stochastic differential equation into a random one. Using this transformation we prove the existence of attractors for the stochastic sine-Gordon equation for a large class of noise terms.

Manthey, Ralf Friedrich-Schiller-Universität Jena, Fakultät für Mathematik und Informatik

Stabilität von Lösungen stochastischer partieller Differentialgleichungen

Im Vortrag werden Ergebnisse zur Stabilität von Lösungen stochastischer partieller Differentialgleichungen mit biologischer oder physikalischer Relevanz vorgestellt.

Maurer, Markus UniBw München, Institut für Wasserwesen

The Effects of Stochastic Input Parameters on Deposision of Suspended Sediment

Calculation of sediment transport processes is a very difficult task. Due to many influencing factors and the impossibibity to develop a fully valid physically model, there exist many approximate models to simulate the process of sediment transport. For engineering predictions in unsteady flow, there exist numerical schemes to solve the St. Venant equations in one or two dimensions coupled with an equation for sediment tranport. The problem is not only the accurate calibration of the model, but also the acquisition of realistic values for the various input parameters. For most of the parameters we have to consider their spatial and/or temporal stochastic nature, and therefore we must describe them as random variables or spatiotemporal random fields. The well known first order reliability method is a useful tool to assess the influence of stochastic input variables to the calculated result. In this paper we will show how this method can be applied to judge quantitatively the influence of the inaccuracy of different input parameters with a simple model for deposition of suspended material in one-dimensional unsteady flow.

Mauthner , Susanne Technische Universität Darmstadt

Schrittweitensteuerung bei der numerischen Lösung stochastischer Differentialgleichungen

Die meisten Integratoren zur Lösung deterministischer Differentialgleichungen enthalten eine Schrittweitensteuerung. Diese bewirkt, daß in Bereichen, in denen die Lösung ”glatt“ ist, eine relativ große Schrittweite gewählt wird, während die Schrittweite in ”nicht-glatten“ Bereichen der Lösung verkleinert wird. Wir stellen einen Ansatz zur Steuerung der Schrittweite bei der numerischen Lösung stochastischer Differentialgleichungen vor: Unter Verwendung der von K. Burrage hergeleiteten Bedingungsgleichungen für die Koeffizienten eines stochastischen Runge-Kutta-Verfahrens konstruieren wir ein eingebettetes Runge-Kutta-Verfahren der Ordnung 1.5(1.0). Dieses eingebettete Runge-Kutta-Verfahren besteht aus einem Runge-Kutta-Verfahren der starken Konvergenzordnung 1.0 und aus einem, auf diesem Verfahren der Ordnung 1.0 aufbauenden, Runge-Kutta-Verfahren der starken Konvergenzordnung 1.5. Während eines der Verfahren zur Approximation der exakten Lösung verwendet wird, dient die Differenz der beiden Verfahren in jedem Integrationsschritt zur lokalen Genauigkeitskontrolle und somit zur Steuerung der Schrittweite. In beiden Verfahren treten neben den Zuwächsen des Wiener-Prozesses noch weitere iterierte Integrale aus der stochastischen Ito-Taylorentwicklung auf. Wir behandeln das Problem ihrer Simulation bei einem Verfahren mit Schrittweitensteuerung. Abschließend vergleichen wir anhand einiger Testbeispiele die Effizienz einer Implementierung mit Schrittweitensteuerung mit der einer Implementierung ohne Schrittweitensteuerung.

Potthoff, Jürgen Universität Mannheim

Strong Solutions of the Zakai equation

It is well–known that the stochastic non–linear filtering problem leads to the Zakai equation, i.e., a stochastic partial differential equation of the following type:

du(t, x) = (Lu + cu)(t, x) dt + ∑_{i=1}^{m} h_{i}u(t, x) dY_{t}^{i},

where L is a second order elliptic differential operator, h_{i} and c are real functions, and Y_{t}^{i} = ∫_{0}^{t} ψ^{i}(s) ds+W_{t}^{i}, with a Brownian motion W_{t}. A differential calculus for random fields is developed, which is combined with the S–transform to derive an explicit strong solution of the above equation. If time permits, applications to non–linear filtering and to mathematical finance will be discussed. The talk is based on a joint paper with F.E. Benth, Th. Deck and G. Våge.

Raupach, Peter Private Universität Witten/Herdecke

On Drift–Free SDE In Dimension One

The stochastic differential equation dX_{t} = b(t, X_{t}) dB_{t} can be solved by time-change of a Wiener process W even if the diffusion coefficient depends explicitly on time. The time-change process (A_{t})_{t≥0} which we seek is obtained as a solution of dA_{t} = b^{2}(t, (W ∘ A)_{t}) dt. Solution A is the quadratic variation of the time-changed Wiener process W ∘ A, which is the solution of the SDE after some slight modification. First a new existence result including degenerate diffusion is presented. The time-change process A constructed here turns out to be pure. As a consequence, the corresponding solution of the SDE has the representation property. We find maximal and minimal time-change solutions A under some weak conditions. The corresponding solutions of the SDE can be interpreted as extremely “hot” or “cold” due to the correspondence of temperature and diffusion. Such extreme solutions are pure, too, and the maximal (resp. minimal) solution of (Z) is unique in law. This fact yields a simplification concerning uniqueness: If uniqueness in law is to be shown, it is only necessary to consider pure solutions in case extreme solutions exist.

Scheutzow , Michael TU Berlin

Ausbreitung von stochastischen Flüssen

Wir betrachten stochastische Flüsse auf euklidischen Räumen. Unter verschiedenen Voraussetzungen an die lokalen Charakteristiken studieren wir das Wachstum des Durchmessers des Bildes einer beschränkten Teilmenge unter dem Fluss. Weiterhin zeigen wir Abschätzungen für die Änderung des Flusses als Funktion des Ortes innerhalb einer festen Zeitspanne. Wesentliche Hilfsmittel sind das Lemma von Garsia-Rodemich-Rumsey und/oder das aus der Theorie der empirischen Prozesse bekannte Chaining. Mit Hilfe von Beispielen demonstrieren wir die Qualität der Abschätzungen. Fragestellungen dieser Art sind zum Beispiel in der Ozeanographie (Ausbreitung von Ölteppichen) von Interesse. Der Vortrag basiert auf gemeinsamen Arbeiten mit M. Cranston, P. Imkeller, S. Mohammed und D. Steinsaltz.

Schurz, Henri La Universidad de Los Andes, BOGOTA

Linear- and Partial-Implicit Numerical Methods for Stochastic Differential Equations

The problem of numerical approximation of high-dimensional stochastic ordinary differential equations (SDEs) arises through discretization of stochastic partial differential equations (SPDEs) in space and time (e.g. by use of stochastic counterpart of method of lines). Then, in particular, very efficient, convergent and asymptotically stable stochastic-numerical methods are needed. The use of fully implicit numerical methods is obviously restricted by local resolution of large, nonlinear and implicit stochastic-algebraic systems. To reduce the total computational effort we shall present more efficient numerical methods, like linear- and partial-implicit methods under the presence of stochasticity. The convergence of corresponding numerical trajectories is proven in the case of SDEs driven by White noise within a stochastic L^{2}-analysis. Thanks to good stability behaviour (see monograph Schurz (Logos-Verlag, Berlin, 1997)), linear- and partial-implicit midpoint methods can be prefered in view of practical implementation. Eventually, we shall apply presented techniques to approximate the paths of stochastic Duffing oscillators (i.e. with terms of cubic nonlinearity) and compare the computational efficiency of them with known stochastic-numerical methods.

References:

H. Schurz: Stability, stationarity, and boundedness of some numerical methods for stochastic differential equations, Logos-Verlag, Berlin, 1997.

Sobiechowski, Carsten Institut für Mechanik, Otto–von–Guericke–Universität Magdeburg

Stochastische Differentialgleichungen für durch Poissonsche Impulse erregte Systeme

Nach einer kurzen Einführung in die Poissonschen Impulsprozesse werden zwei Konzepte zur Beschreibung von Systemen unter Poissonscher Impulsanregung vorgestellt, wobei insbesondere der Anwendbarkeit auf ingenieurmäßige Probleme Beachtung geschenkt wird. Im zweiten Teil des Vortrages stehen dann stationäre Wahrscheinlichkeitsdichtefunktionen für den Systemausgang im Mittelpunkt. Es wird die Entsprechung der Fokker–Planck–Gleichung angegeben, die dann für externe oder linear parametrische Impulserregung mit deterministischen oder n–Punkt verteilten Amplituden spezifiziert wird. Durch Vorgabe von Klassen von Wahrscheinlichkeitsdichtefunktionen lassen sich dann Systeme ermitteln, welche diese Funktionen als stationäre Wahrscheinlichkeitsdichtefunktionen annnehmen. Die Klassen von Systemen, für die exakte stationäre Wahrscheinlichkeitsdichtefunktionen ermittelt werden können, enthalten insbesondere die von Vasta [1] angegebenen Klassen als Spezialfall. [1] Vasta, M.: Exact stationary solution for a class of non–linear systems driven by a non–normal delta–correlated process, International Journal of Non–Linear Mechanics 30 (4), p. 407–418, 1995

Trutnau , Gerald Fakultät für Mathematik, Universität Bielefeld

Stochastic calculus of generalized Dirichlet forms

We show Fukushima’s decomposition of AF’s in the frame work of quasi-regular generalized semi-Dirichlet forms considered in W. Stannat, Generalized Dirichlet forms and associated Markov processes, C.R. Acad. Sci. 319, 1063-1069 (1994). The only condition assumed, is that the co-resolvent is sub–Markovian. No dual process is needed. The key-point for the proof of the decomposition is an integral representation theorem for coexcessive functions and a resulting description of Eexceptional sets by an appropriate class of measures. We also derive an Itô-type formula for the transformation of the martingale part of the decomposition.

Zähle , Martina Mathematisches Institut der Universität Jena

On the link between fractional and stochastic calculus

Methods of classical fractional calculus are applied to generalized Stieltjes and stochastic integration theory. For processes with generalized covariations this may also be interpreted in terms of an extension of a concept of Russo and Vallois. Finally we consider SDE under these aspects.

Antoniouk, Alexander Val. & Alexandra Vict. Kiev Institute of Mathematics, Acad.Sci.Ukraine

Nonlinear Calculus of Variations in Applications to the C^{∞}-Properties of General non-Lipschitz Diffusions

In this short lecture we develop further a program on study the stochastic differential equations with essentially nonlinear diffusion and drift coefficients via the observation on the nonlinear symmetries of high order variational calculi. Such symmetries lead to the nonlinear a priori estimates on quasi-contractive in time behaviour of variations and permit the investigation of C^{∞} properties of associated Feller semigroups. The complete scheme in simple IR^{d}-valued coercitive diffusion case is discussed and we demonstrate the influence of non-Lipschitzness parameters on the differentiability of solutions with respect to initial data and Wiener space variables. The applications to the C^{∞} smooth and smoothing properties of discontinuous Feller semigroups are obtained. We also derive the high order generalizations of Elworthy formula for the heat semigroups derivatives and point on a new effect that one should introduce a non-Lipschitz gap between boundedness and continuity in the topologizations of C^{n} spaces, preserved by transitional semigroup. The infinite dimensional extensions of results are briefly sketched.

Key words: Nonlinear SDE, finite and infinite dimensional Feller semigroups, C^{∞} differentiability of solutions, nonlinear estimates on variations, raise of smoothness and Elworthy formula via Malliavin calculus.